wiki:UC20171211

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 12/11/17

Attending: Nancy Chanover (NMSU; minute taker), Mark Klaene (APO), Rachel Kuzio de Naray (GSU), Joanne Hughes Clark (Seattle U), Kal Kadlec (UW), Sarah Tuttle (UW), Mary Beth Kaiser (JHU), Bill Ketzeback (APO), Scott Anderson (UW), John Bally (CU), Ben Williams (UW), Mark Whittle (UVa)


User feedback and comments from institutional representatives:

  • Wyoming - no report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - nothing to report
  • Oklahoma - no report
  • NMSU - good class trip, scattered light reported from DIS data reduced by student, NC will pass on to Russet, Bill, Mark, Sarah, Kal
  • NAPG - no report
  • JHU - observing run on Sat night - lost 1.5 hrs due to pointing issue
  • Georgia State - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 11/09/17 through 12/05/17

0) Overview

Reasonably good weather this period. ARCSAT observing. Signs of winter are just starting to show. DIS and Echelle scattered light being monitored closely; both are holding steady but at values worse than we thought were good.

1) Telescope

Telescope was operational with minimal issues. We had one issue with tertiary actuator that failed to provide encoder values due to a bad connector caused by getting snagged during rotation. Bill and Ed came in on Sunday to repair.

2) Instruments

DIS: The scattered light trend is still generally worse than we like. Winter temperatures will severely limit any servicing over the next few months.

TripleSpec is operational.

Agile has been operational this period with no issues.

Echelle is operational. Inter-Order Light is higher than we like but improving slightly. Ecam shutter (new spare installed a couple weeks ago) is causing some issues. Bill serviced it now a couple times with some improvement. New shutters are no longer available except in large lots and the original shutter is out for refurbishment. Estimate return date is TBD from manufacturer. There is also readout banding that is causing some issues. A replacement camera order is in the works as Andor repair service has gotten terrible.

NIC-FPS is operational.

ARCTIC is operational.

3) ARCSAT - General operations have been nominal with some issues on scripting hangups and other computer/software glitches. SurveyCam is still at Andor with no ETA.


Additional telescope and instrument discussion:

Instruments: we are still monitoring scattered light in both DIS and ARCES. It is harder to service instruments during the winter months due to the colder temperatures in the dome; barring any major problem those two instruments won’t be serviced in the very near term. TripleSpec, Agile, ARCTIC, NICFPS have all been good, with a slight improvements in the inter order light for ARCES - this is possibly due to a better ion pump or colder temperatures. Ecam has been troublesome with ongoing readout issues and shutter problems. Bill serviced the shutter (so the spare is in there now). Replacing the shutter is challenging because Melles Griot now requires that shutters be purchased in very large lots so we sent back the original for repair. Spare boards were installed but the problem came back. Due to unsatisfactory service from Andor we are now moving forward with the purchase of a Finger Lakes camera that will become the new Ecam (the current one will become the spare). The new camera will have to be integrated into TUI, ICC, guiding software, etc., so it will require some software development.

ARCSAT: we received word that the SurveyCam work is complete at Andor and the camera is now on its way back to us. In general ARCSAT is operating nominally, and we are making plans to move forward with getting M1 resurfaced and aluminized with an enhanced overcoat. It will be offline for most of Q1 and will return to service in the spring.

Pointing issue: On Thursday and Friday nights there was evidence that the pointing had changed dramatically on the telescope. Science was preempted on Friday night to do a small pointing model, which indicated that dramatic changes to the way the telescope is leveled to the sky had shifted. The engineering staff is planning to preempt science tonight to further evaluate it and correct it. They have not seen this happen as dramatically in the past; they have seen shifts in pointing due to thermal changes but not like this, or not in this way. They are trying to address the issue as quickly as possible. Mary Beth asked whether the earlier crude pointing model from Saturday night held or not. It held (sort of) for instruments at NA2 - it was OK for DIS but once they switched over to the echelle they struggled to do a pointing correction using the mirrors (primarily tertiary) of the telescope rather than local corrections on sky. It took lots of effort and time to get the echelle pointing good enough in order to observe during B half.


Q4 3.5-m scheduling:

All remaining OPEN/DD time has been allocated; the schedule is complete.


Q4 ARCSAT scheduling:

No changes.


Q1 3.5-m scheduling:

The schedule should be out in a few days. There were a few conflicts but they have been resolved already, no major complications for Q1 2018.


Q1 ARCSAT scheduling:

There was no call for Q1 but there is a possibility of getting a little time in early January.


TDAWG Update

The Time Domain Astronomy Working Group (TDAWG) is going strong and holding regular telecons, and now has a solid draft of a report put together. They are currently working to learn more about what the different institutions are interested in. Ben sent an email to the institutional schedulers and so far only received one response - he requested that schedulers poll their users and also talk to their UC reps. The idea is to gauge whether people are willing to be preempted (this can already be indicated on the proposal cover sheet) all the way to whether people are interested in cross institutional collaborations to pool time and dedicate it for time domain follow up. Ben suggests that the UC reps make a point of discussing this with their institutional schedulers as well as their department users to get a sense of what people in their departments are interested in.


TripleSpec Slit Width

There were no comments from the UC reps concerning science drivers for observing with alternate slit widths (i.e. not the default of 1.1"). We could also consider defining a single (but different) default slit width if people desired that, but so far Nancy has received no input indicating that anyone wants to use anything aside from the 1.1" slit.


NA2 Baffle Project

Several members of the UVa team have been unavailable for the last several weeks so there was not much update. Their team is still gathering information to develop a SolidWorks model that could eventually be imported into Zemax. We expect to see more on this in early January. ARCTIC flats that are intended to assess the degree of vignetting have been transferred to Nancy. She will make the link for the repository available on the ARCTIC page of the users wiki.


Strategic Planning Exercise in 2018 (preliminary discussion)

Nancy provided the UC with a brief overview of her plans for a strategic planning session in 2018 to develop a longer-term vision and plans for the telescope. This will aid in informing decisions concerning instrumentation, scheduling, partnerships, the use of the telescope for educational purposes, etc. The goal is to have this effort completed in time for Nancy to present it to the Board of Governors in November 2018. This discussion was simply to notify the UC that this is in the works, and that people will be hearing more about it soon.


ACTION ITEMS:

open from previous meetings:

  • Russet will iterate with the UVa team to clarify some of the language in the TripleSpec User's Guide. Status: OPEN. Still ongoing. Russet needs to narrow down specific questions that are not clear from the manual, then Mike S and John W will help clarify the language.
  • Ben and Nancy will discuss the ARCTIC flats issue on an upcoming 3.5m engineering telecon to figure out what is acceptable for the Observing Specialists, and come up with a policy regarding how late the observers can keep them to take flats. Status: OPEN. Nancy compiling a policy that might go on users' wiki.
  • For the NA2 baffle project, Joanne will send a flat field image around to work with. Based on the information from Joanne Nancy will work with John Wilson to add the amplitude of the current loss of signal in the corners of the ARCTIC FOV to the NA2 Baffle Project Description document. Status: OPEN. All of the relevant files were transferred to Nancy. She will post a link to them on the ARCTIC users wiki page.
  • All UC reps: Please notify users of constraints on Tspec slit use and forward any interest in using other slits to committee. Status: CLOSED. No input received so we will leave TripleSpec at the default 1.1" slit width until we hear otherwise.

new from this meeting:

  • All UC reps: coordinate with your institutional scheduler to get a sense from your users concerning the degree to which they would be willing to be preempted for time domain followup work, and/or whether they would be interested in pooling time across ARC institutions to dedicate to time domain followup observations.

Next meeting:

The next Users Committee telecon will be on Monday February 5 at 11 am MST. Nancy will send out another query to all UC reps to confirm that this time is still acceptable in the spring. We will reserve Jan. 22 for an as-needed telecon if there are any late-breaking issues that need to go before the Users Committee.

Note: See TracWiki for help on using the wiki.